# A Goemans-Williamson type algorithm for identifying subcohorts in clinical trials

### v1

@@ Abstract
We design an efficient algorithm that outputs a linear classifier for identifying homogeneous subsets (equivalently subcohorts) from large inhomogeneous datasets. Our theoretical contribution is a rounding technique, similar to that of Goemans and Williamson (1994), that approximates the optimal solution of the underlying optimization problem within a factor of $0.82$. As an application, we use our algorithm to design a simple test that can identify homogeneous subcohorts of patients, that are mainly comprised of metastatic cases, from the RNA microarray dataset for breast cancer by Curtis et al. (2012). Furthermore, we also use the test output by the algorithm to systematically identify subcohorts of patients in which statistically significant changes in methylation levels of tumor suppressor genes co-occur with statistically significant changes in nuclear receptor expression. Identifying such homogeneous subcohorts of patients can be useful for the discovery of disease pathways and therapeutics, specific to the subcohort.
@@ 1 Introduction
A linear separator in $\mathbb{R}^{d}$ is a basic tool used in statistics and machine learning to partition a set of points into two subsets. Such a separator implicitly encodes information about the pairs of points, which it has to separate, in its coefficients. Given a set of Red and Blue points to be separated, there may not exist a linear classifier that separates most of the Red and Blue points. Moreover, even if one exists, then it can be expected to have most coefficients to be far from zero, i.e., it’s not sparse, unless there are structural assumptions on the dataset. On the other hand, if we want a linear separator that separates a small (but non-trivial) percentage of Red points from Blue points, then we may reasonably expect that a sparse linear separator exists, i.e., one which sets most of its coefficients to zero, and it still has low error. This is because the number of Red-Blue pairs separated is smaller, so the amount of information to be encoded in the linear separator is lower. Such separators also have the advantage that they are explainable – the role and importance of each feature is clear from its coefficient. Therefore, if one assumes the separator is sparse, the focus shifts to identifying a subset of Red points to carve out from the rest, so that: (1) the number of Blue points in the identified subset is relatively small (specificity is high); (2) the relative number of Red points in the subset, compared to the total number of Red points, is large (sensitivity is high); and (3) the number of non-zero coefficients in the separator is small (the separator is sparse). Computing such linear separators are the subject of this paper, and we devise a semidefinite programming (SDP) based algorithm (see Algorithm 1) for that purpose. Furthermore, we prove theoretical guarantees on the size of the identified subset (see Theorem 2.2).
One application for such an algorithm is in identifying homogeneous subsets (equivalently subcohorts) of patients in clinical trials. Diseases like breast cancer, sepsis and even bacterial infections are fairly inhomogeneous – the same disease may have very different underlying causes (different mutations, different strains of bacteria, and so on), and therefore different treatments. No single treatment works for a majority of the patients. Therefore, given a drug that works on a small percentage of patients overall, one recourse is to identify a not too small subset for which the drug treatment works with higher likelihood. Moreover, one wants the identifying test to rely on a handful of features (few genes, few mutations etc) – to ensure explainability of the test, and also control for sequencing costs in the field. This is precisely the abstract problem delineated in the previous paragraph – we are willing to trade sensitivity for specificity. Thus, in this paper, we use the algorithm mentioned above (Algorithm 1), on the RNA microarray dataset for breast cancer from CSC+ (12); PCR+ (16); RSS+ (19) [^1] We use this particular dataset because it is large and balanced, i.e., both metastatic and non-metastatic cases are adequately represented. Moreover, it has epigenetic data as well as RNA expression data, and linking the two is an interesting open question in biology CH (19). to identify subsets of metastatic breast cancer patients that have statistically significant over-expression of one or more nuclear receptors, compared to the full dataset (see Figure 3). It’s worth noting here that nuclear receptors are often chosen as drug targets OALH (06).
@@ 1.1 Related literature
A two-class linear classifier seeks to separate all Red-Blue points in the input, thus the focus is on minimizing the mean-square error, which is a tractable problem. In our problem, we assume that either a low error linear classifier does not exist, or even if it exists, it is not supported over a small set of co-ordinates. Thus, the challenge in our problem is to identify the correct subset. It resembles a version of the MAX-CUT problem GW (95) – we want to identify a linear classifier that cuts as many Red-Blue edges as possible, and ensures that the side with Red points has relatively few Blue points. Moreover, we want most of the coefficients of classifier to be close to zero. This class of problems are not as tractable – most versions of the problem are computationally hard (cf. balanced cut, sparsest cut ARV (09) and so on). However, it turns out that efficient (similar in efficiency to computing an optimal linear classifier) approximate algorithms can and do exist. Moreover, the approximation guarantee of some such algorithms is pretty close to $1$ (for e.g. GW (95)). In Subsection 2.1, we show that formulating our problem as a 00-$1$ optimization problem (similar to MAX-CUT), and using a SDP based rounding algorithm, allows us to achieve similar efficiency as a linear classifier, and also obtain desired sensitivity vs specificity trade-offs.
Finally, while algorithms for cut problems, like balanced cut, have been used for clustering BLUvB (13), they are not specifically designed to account for inputs embedded in $\mathbb{R}^{d}$. On the other hand, one can think of the identified subcohort as a single homogeneous cluster, and potentially use a standard clustering algorithm operating on inputs in $\mathbb{R}^{d}$ (like Lloyd’s algorithm or other $k$-means algorithms). However, some key difficulties will remain. First, such algorithms compute non-linear separators, which can lead to further loss of explainability. It is unlikely that one can control the solution sparsity [^2] This would be the number of genes used in the test in our application. for such non-linear separators. For example, in the case of Lloyd’s algorithm, the subset of points for the identified subcohort would be bounded by the corresponding Voronoi polytope, which may have an exponential number of facets in the dimension $d$ – leading to a non-linear solution that’s also unlikely to be sparse. Secondly, $k$-means clustering algorithms can be computationally intractable as well – versions of $k$-means that achieve approximation guarantees close to $1+\epsilon$, have time complexity super-exponential in $\frac{1}{\epsilon}$ (see for example KSS (04)). Thirdly, it is unclear how to set the parameters to control the sensitivity and specificity when using a cookie-cutter clustering algorithm for $k$-means. While the size of clusters (proportional to the sensitivity) could be controlled using constraints, it is unclear how to optimize for specificity explicitly in known constrained $k$-means implementations. Finally, while explainable clustering has been studied for $k$-means type algorithms before (for e.g. MDRF (20) and GPSY (23)), such algorithms are restricted to axis-aligned thresholding cuts, as opposed to general but sparse cuts (as in this paper). Thus their techniques and results are significantly different from the ones outlined in this paper.
@@ 2 Overview of our results
We want to solve the following high-level problem. Suppose one is given a set of patients which have the same disease but different disease states. For example, in case of cancer patients, some of them may have metastasis but others may not. We want to identify a small set of genes or proteins (or some other markers) that allows us to test and select patients that are in the same state, i.e., we want to design tests that identify predominantly homogeneous subcohorts (subsets) of patients. For example, if roughly 10% of the patient population has metastasis, then we want to design an algorithm that uses a small number of markers to identify a large enough subcohort, say 2% of the overall patient population, of which 70% have metastasis.
@@
Clearly, there are trade-offs involved. If we want the identified subcohort of patients to be a larger fraction of the overall patient population, i.e., if we want the sensitivity to be large, then the probability that most patients in the subcohort will have metastasis, will be smaller, i.e., the specificity will be smaller (equivalently the subcohort will be more inhomogeneous). Furthermore, if we want the set of identifying markers to be small, i.e., if we want the designed test to be sparse, then the specificity is going to be lower (assuming sensitivity is kept constant). On the other hand, if we want to have sparse tests and high specificity in our subcohort, then the sensitivity (measured by subcohort size) will be reduced.
@@
In this paper, we formalize and solve the subcohort identification problem using Algorithm 1,which is a generalization of the Goemans-Williamson semidefinite programming (SDP) algorithm for solving MAX-CUT GW (95). Our theoretical contribution can be summarized as follows:
@@
  • We prove that the ratio of the size of subcohorts that are identified by Algorithm 1 vs. the optimal subcohorts is lower bounded by an absolute constant $\alpha\simeq 0.82$ (see Theorem 2.2 and Subsection 2.1 for details).
@@
The theoretical result should be independently interesting, as it generalizes the Goemans-Williamson calculations. While SDP rounding algorithms for similar problems have been analyzed before (for e.g. FG (95); LLZ (02) and BHPZ (20)), we are not aware of any similar analysis for the problem in our paper. [^3] In a nutshell, the crux of the Goemans-Williamson guarantee relies on computing the maximum ratio of the chord length to the minor arc length in a unit circle, while in our case it hinges on bounding the ratio of the ”perimeter” of a triangle to the perimeter of the circumcircle, the latter is assumed to be of unit radius. It is surprising that a Goemans-Williamson SDP rounding works here, as our objective function is not sign definite. Usually more complicated rounding techniques, like AN (06) are required in such cases.
@@
Furthermore, we empirically study the trade-offs mentioned above, for Algorithm 1, in the context of identifying subcohorts of metastatic breast cancer using the METABRIC dataset CSC+ (12); PCR+ (16); RSS+ (19). [^4] The dataset contains RNA microarray data of approximately 2,000 breast cancer patients, and methylation data for approximately 13K genes in about 1,500 breast cancer patients. About a thousand patients had one or more positive lymph nodes upon biopsy (metastasis), while the remaining did not (non-metastasis).
@@
For the empirical part, our goal was to identify subcohorts of predominantly metastatic breast cancer cases using tests based on expression of nuclear receptors, i.e., we want the ratio of the metastatic cases to non-metastatic cases to be high for the identified subcohort. We denote this ratio as the specificity of our test, the percentage of metastatic cases in the identified subcohort is denoted as the sensitivity of the test. The number of genes used is roughly proportional to the sparsity parameter of Algorithm 1. Our first empirical result relates the specificity, sensitivity and the number of genes used, as follows:
@@
  • We show that specificity and sensitivity both increase with the number of genes used in the test (see Figure 1(a)). In particular, we obtained subcohorts of size $\sim 20$% of the subsampled patient population, with about 57% specificity, when we use roughly 15 out of 48 nuclear receptors in the designed test.
@@
  • Moreover, if we define our subcohort as an intersection of two subcohorts, from two different runs of Algorithm 1 (over two subsamples of the dataset), then we can increase the specificity by sacrificing on sensitivity (see Figure 1(b)). In particular, we obtained subcohorts of size $\sim 7$% of the subsampled patient population, with about 67% specificity, when we use roughly 45 out of 48 nuclear receptors in the test.
@@
Our second empirical result, is an application of our algorithm to identify statistically significant associations between sets of genes. In this case, we identify statistically significant associations between epigenetic data and downstream nuclear receptor expressions. For example, in the context of the breast cancer dataset, we use subcohorts identified using Algorithm 1 to associate statistically significant changes in methylation of certain known breast cancer related genes to statistically significant changes in expression of certain nuclear receptors (see Figure 3). Identifying statistically significant associations between changes in methylation percentages and changes in nuclear receptor expressions is an interesting problem since methylation of certain genes is known to be a potential cause of progression of breast cancers SK (15), while nuclear receptors are often used as targets of many drugs BSW+ (13). Therefore, we are attempting to link the likely cause of the pathology to a likely treatment target, specific to a subcohort of patients.
@@
The question arises: why is identifying such associations difficult or novel, after all, it involves only correlation computation? A naive average over the entire cohort of patient population, where one indiscriminately averages over non-homogeneous data, may make confidence intervals larger than the separation between the means. Thus it becomes difficult to say whether any identified associations are statistically significant. This is exactly what happens in Figure 2, where we compute the average expression of nuclear receptors and methylation percentages in metastatic and non-metastatic breast cancer patients, over the entire cohort of about 2,000 patients. Based on the figure, one can not differentiate between metastatic vs non-metastatic breast cancer using correlations and simple z-scores of any of the 90 odd genes plotted – the confidence intervals are not narrow enough to separate the means. Similarly, if one computes averages over arbitrary small subsets of the population, across subcohorts, then the confidence intervals are likely to be too large as well. Therefore, we need to average over "good" (homogeneous) subcohorts, i.e., subcohorts where averaging likely leads to narrower confidence intervals but not closely clustered gene expressions (see Subsection 2.2.2 for details). Identifying subcohorts based on a small set of markers should implicitly help ensure homogeneity within the subcohort, and narrow confidence intervals. Using subcohorts identified by Algorithm 1, we obtain the following result for the breast cancer dataset:
  • We used Algorithm 1 to generate several thousands of subcohorts by subsampling the input (one subcohort for each subsampled input). One identified subcohort, together with the resulting statistically significant association between methylation percentages of known breast cancer related genes and nuclear receptor expressions, is shown in Figure 3. For the subcohort in Figure 3, the significant over-expression of LXRB receptor gene suggests an initial hypothesis: patients in the subcohort may be good candidates for LXR-inverse agonists, if supported by clinical evidence.
@@
    Here, it is worthwhile to note that both LXR-inverse agonists CVS+ (19) and LXR agonists HLRL+ (19) have been proposed to be beneficial for various subsets of breast cancer patients. Our algorithm, and its output test, may offer a principled way to identify genomic pathways underlying the disease. Moreover, it may also help identify patients, who could benefit from use of LXR-inverse agonists or LXR agonists (or neither).
Furthermore, in actual practice, the physician may expect to be given genetic sequencing data from a single patient, so it is natural to ask: Is there a way to decide if a given sequenced sample belongs to some "good" subcohort? If it does, the physician may be able to use the computed statistically significant associations between the genes in that subcohort to arrive at an appropriate treatment. So the answer to the question above is likely to be affirmative, as long as one can build a large library that matches a given genetic sequence to that from a precomputed homogeneous subcohort. It’s worth noting here that there are potentially exponentially many subcohorts that one may need to generate, and then identify good subchohorts from amongst them, in order to build a library of homogeneous subcohorts. Thus, one may need to train a neural net to prioritize identification of subcohorts that are more likely to lead to statistically significant associations. We found that the problem is very amenable to learning using neural nets, and for the breast cancer dataset, even a single hidden layer neural net has reasonable accuracy (see Figure 4 and Subsection 2.2.3 for details).
@@
We begin with formal algorithmic description and theoretical details in Subsection 2.1, which is followed by empirical results on the breast cancer dataset in Subsection 2.2. The biologically minded reader may skip Subsection 2.1, while the mathematically minded reader may skip Subsection 2.2, upon first reading.
2.1 Underlying theoretical problem

### v2

@@ Abstract
We design an efficient algorithm that outputs tests for identifying predominantly homogeneous subcohorts of patients from large in-homogeneous datasets. Our theoretical contribution is a rounding technique, similar to that of Goemans and Wiliamson (1995), that approximates the optimal solution within a factor of $0.82$. As an application, we use our algorithm to trade-off sensitivity for specificity to systematically identify clinically interesting homogeneous subcohorts of patients in the RNA microarray dataset for breast cancer from Curtis et al. (2012). One such clinically interesting subcohort suggests a link between LXR over-expression and BRCA2 and MSH6 methylation levels for patients in that subcohort.
@@ 1 Introduction
A linear separator in $\mathbb{R}^{d}$ is a basic tool used in statistics and machine learning to partition a set of points into two subsets. Given a set of Red and Blue points to be separated, there may not exist a linear separator that separates most of the Red and Blue points. Moreover, even if one exists, then it can be expected to have most coefficients to be far from zero, i.e., it’s not sparse, unless there are structural assumptions on the dataset. On the other hand, if we want a linear separator that separates a small (but non-trivial) percentage of Red points from Blue points, then a sparse linear separator may exist. Such sparse separators also have the advantage that they are likely to be interpretable. Therefore, the focus shifts to identifying a subset of Red points to carve out from the rest, so that: (1) the number of Blue points in the identified subset is relatively small (specificity is high); (2) the relative number of Red points in the subset, compared to the total number of Red points, is large (sensitivity is high); and (3) the number of non-zero coefficients in the separator is small (the separator is sparse).
In this paper, we devise a semidefinite programming (SDP) based algorithm (see Algorithm 1) for computing sparse linear separators that identify homogeneous subgroups. We prove theoretical guarantees on the specificity and sensitivity of the separator (see Theorems 3.2 and 3.3). We apply our algorithm to design a test for identifying subgroups of metastatic breast cancer cases from a large breast cancer dataset metabric; metabric2; metabric3. We empirically study the trade-offs between the sensitivity, specificity and sparsity of the tests output by our algorithm in Section 4.1. Finally, in Section 4.2, we use our algorithm to identify clinically interesting subgroups of breast cancer patients.
@@ 1.1 Related literature
Perhaps the most well-known subgroup identification algorithm is the Patient Rule Induction Method (PRIM) ff and its variants rlbk; prim-cart. These are local search algorithms with heuristics designed to gradually trade-off sensitivity for specificity. At it’s crux, the better-known PRIM algorithm relies on identifying one or more hyper-rectangles, where points in the interior of a hyper-rectangle constitute the identified subgroup. While such bump hunting heuristics are hard to analyze theoretically, explainable clustering algorithms using axis-aligned thresholding cuts have been theoretically analyzed (for example SIELING2008394expl-kmeans1 and expl-kmeans2). There are at least two significant differences between algorithms using axis aligned hyper-rectangles (equivalent to thresholding rules on coordinates) and our approach. First, such rule-based algorithms are almost always restricted to axis-aligned thresholding cuts, as opposed to sparse cuts (as in this paper); even though the latter are also intuitively interpretable. Second, each axis aligned half-space can also be thought of as node in a decision tree, so computing subgroups that are identified by one or more axis-aligned hyper-rectangles corresponds to computing a decision tree, where the underlying objective is chosen to balance sensitivity and specificity. It is well-known that computing optimal decision trees for most reasonable objectives is not just a NP-hard problem, but most versions are even hard to approximate within any constant factor (see for e.g. gupta1; koch2024superconstant). Intuitively, this is because of close connections between decision trees and circuit complexity. In contrast, in this paper, we translate subgroup identification with sparse linear cuts to a MAX-CUT like problem (cf. gw), which admits constant factor approximation guarantees via convex optimization (an approximation factor of $0.82$ in our case). Moreover, there are practical examples where a sparse linear hyperplane, as opposed to a rule based algorithm, is used to identify interesting subcohorts. For example, the papers fbtest and meld describe diagnostic scores, essentially sparse regression models, that are widely used in gastroenterology.
@@ 2 Overview of our results
Motivation: Suppose we are given a set of patients in different disease states. For example, in case of cancer patients, some of them may have metastasis but others may not. We want to identify a small set of genes or proteins (or some other markers) that allows us to test and select patients that are in the same state, i.e., we want to design tests that identify predominantly homogeneous subcohorts (subsets) of breast cancer patients with metastasis. For example, if roughly 10% of the cancer patients have metastasis, then we want to design an algorithm that uses a small number of gene expressions to identify a large enough subcohort, say 2% of the overall patient population, of which 70% have metastasis.
@@
Clearly, there are trade-offs involved. If we want the identified subcohort of patients to be a larger fraction of the overall patient population, i.e., if we want the sensitivity to be large, then the probability that most patients in the subcohort will have metastasis, will be smaller, i.e., the specificity will be smaller (equivalently the subcohort will be more inhomogeneous). Furthermore, if we want the set of identifying markers to be small, i.e., if we want the designed test to be sparse, then the specificity is going to be lower (assuming sensitivity is kept constant). On the other hand, if we want to have sparse tests and high specificity in our subcohort, then the sensitivity (which corresponds to subcohort size) will be reduced.
@@
Contribution: In this paper, we formalize and solve the subcohort identification problem using Algorithm 1, which is a generalization of the Goemans-Williamson semidefinite programming (SDP) algorithm for solving MAX-CUT gw. Our theoretical contribution can be summarized as follows:
@@
  • We prove that the ratio of the size of subcohorts that are identified by Algorithm 1 vs. the optimal subcohorts is lower bounded by an absolute constant $\alpha\simeq 0.82$ (see Section 3 for details). This implies bounds on the sensitivity and specificity as well (Theorems 3.2 and 3.3).
@@
The theoretical result should be independently interesting, as it generalizes the Goemans-Williamson calculations. While similar SDP rounding algorithms for similar problems have been analyzed before (for e.g. feige1; rpr2 and potechin), we are not aware of any similar analysis for the problem in our paper. [^1] In a nutshell, the crux of the Goemans-Williamson guarantee relies on computing the maximum ratio of the chord length to the minor arc length in a unit circle, while in our case it hinges on bounding the ratio of the ”perimeter” of a triangle to the perimeter of the circumcircle, the latter is assumed to be of unit radius. It is surprising that a Goemans-Williamson SDP rounding works here, as our objective function is not sign definite. Usually more complicated rounding techniques, like alon-naor are required in such cases.
@@
Furthermore, we empirically study the trade-offs mentioned above, for Algorithm 1, in the context of identifying subcohorts of metastatic breast cancer using the METABRIC dataset metabric; metabric2; metabric3.222The dataset contains RNA microarray data of approximately 2,000 breast cancer patients, and methylation data for approximately 13K genes in about 1,500 breast cancer patients. About a thousand patients had one or more positive lymph nodes upon biopsy (metastasis), while the remaining did not (non-metastasis).
@@
For the empirical part, our first result identifies subcohorts of predominantly metastatic breast cancer cases using sparse linear tests based on expression of nuclear receptors, i.e., we want the ratio of the metastatic cases to all breast cancer cases to be high for the identified subcohort. We denote this ratio as the specificity of our test. The number of metastatic cases in the identified subcohort as a percentage of all breast cancer cases is denoted as the sensitivity of the test. The number of genes used is roughly proportional to the sparsity parameter of Algorithm 1. Our first empirical result relates the specificity, sensitivity and the number of genes used, as follows:
@@
  • We show that specificity and sensitivity both increase with the number of genes used in the test (see Figure 1(a) and Subsection 4.1 for details). In particular, we obtained subcohorts of size $\sim 20$% of the subsampled patient population, with about 57% specificity, when we use roughly 15 out of 48 nuclear receptors in the designed test.
@@
  • Moreover, if we define our subcohort as an intersection of half-spaces from two different runs of Algorithm 1 (over two subsamples of the dataset), then we can increase the specificity by sacrificing on sensitivity (see Figure 1(b)). In particular, we obtained subcohorts of size $\sim 7$% of the subsampled patient population, with about 67% specificity, when we use roughly 45 out of 48 nuclear receptors in the test.
@@
Our second empirical result, is an application of Algorithm 1 to identify clinically interesting large homogeneous subcohorts of patients. One example of a clinically interesting problem is to identify statistically significant associations between changes in methylation percentages and changes in nuclear receptor expressions. Methylation of tumor suppressor genes is known to be a potential cause of progression of breast cancers brcgenes, while nuclear receptors are often used as targets of many drugs nrdrugs. Therefore, we are attempting to associate the likely cause of the pathology to a likely treatment target, for a specific subcohort of patients.
@@
The question arises: why is identifying such associations difficult or novel, after all, it involves only standard deviation computation? A naive average over the entire cohort of patient population, where one indiscriminately averages over non-homogeneous data, may make confidence intervals larger than the separation between the means. Thus deviations from the mean are no longer statistically significant. This is exactly what happens in Figure 2, where we compute the average expression of nuclear receptors and methylation percentages in metastatic and non-metastatic breast cancer patients, over the entire cohort of about 2,000 patients. Based on the figure, one can not differentiate between metastatic vs non-metastatic breast cancer using deviations and simple Z-scores of any of the 90 odd genes plotted – the population is inhomogeneous – confidence intervals are too wide to differentiate between the two classes of patients with statistical significance. Similarly, if one computes averages over many small subsets of the population, across subcohorts, then the confidence intervals are likely to be too large due to small sample size. Therefore, to deduce clinically interesting conclusions, we need to identify and compute the deviations over homogeneous subcohorts.
In Subsection 4.2, we trade-off sensitivity for specificity using Algorithm 1 to make the identified subcohorts more homogeneous. One clinically interesting subcohort is shown in Figure 3. For this subcohort, we observe:
  • There is significant methylation increase in genes BRCA2 and MSH6, and nuclear receptor NR1H2 (equivalently LXRB) is significantly over-expressed, both relative to the baseline population. Therefore, one may hypothesize that methylation of the former leads to downstream over-expression of the latter.
  • Furthermore, significant over-expression of LXRB receptor gene suggests that patients in the subcohort may be good candidates for LXR-inverse agonists, if supported by clinical evidence.
@@
    Here, it is worthwhile to note that both LXR-inverse agonists lxr-ant and LXR agonists lxrbeta have been proposed to be beneficial for various subsets of breast cancer patients. Our algorithm, and its output test, may offer a principled way to identify genomic pathways underlying the disease. Moreover, it may also help identify patients, who could benefit from use of LXR-inverse agonists or LXR agonists (or neither).
Currently, clinically homogeneous subgroups of breast cancer patients are often identified by single genes. For example, ER+ patients, i.e., those breast cancer patients with ER nuclear receptor expression above a certain threshold, are treated with a certain protocol and vice versa bc-protocol. This is a likely suboptimal determination of patient subcohorts and resulting treatment protocols. In contrast, our algorithm presents a systematic approach with a potentially better handle on sensitivity, specificity and sparsity trade-offs.
@@
Organization: We begin with formal algorithmic description and theoretical details in Section 3, which is followed by empirical results on the breast cancer dataset in Section 4.
3 Underlying theoretical problem